# Running in circles: practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

### v2

@@ Abstract
Post-clustering inference in scRNA-seq analysis presents significant challenges in controlling Type I error during Differential Expression Analysis. Data fission, a promising approach, aims to split the data into two new independent parts, but relies on strong parametric assumptions of non-mixture distributions, which are violated in clustered data. We show that applying data fission to these mixtures requires knowledge of the clustering structure to accurately estimate component-specific scale parameters. These estimates are critical for ensuring decomposition and independence. We theoretically quantify the direct impact of the bias in estimating this scales parameters on the inflation of the Type I error rate, caused by a deviation from the independence. Since component structures are unknown in practice, we propose a heteroscedastic model with non-parametric estimators for individual scale parameters. This model uses proximity between observations to capture the effect of the underlying mixture on data dispersion. While this approach works well when clusters are well-separated, it introduces bias when separation is weak, highlighting the difficulty of applying data fission in real-world scenarios with unknown degrees of separation.
Keywords: Unsupervised learning, Mixture Model, Post-clustering inference, Type I error, Non-parametric estimation, local variance
@@ 1 Introduction
Clustering encompasses all unsupervised statistical methods that group observations into homogeneous and separated clusters. Widely used in various application fields such as biology or genomics (Jaeger and Banks,, 2023), clustering plays a significant role to uncover or summarize signals contained in different kind of multivariate data. In the context of single-cell RNA-seq (scRNA-seq) technologies which are high-throughput genomics techniques that measure gene expression at the single-cell level, providing insights into cellular heterogeneity and functional diversity within complex biological tissues, cluster analysis is the first step of the traditional pipeline for data analysis (Amezquita et al.,, 2020). The clustering groups cells based on their gene expression/abundance. Afterwards, the differential expression analysis, comparing gene abundance between groups, allows to identify and annotate cellular sub-populations. This leads to the identification of marker genes that could potentially serve as cell-type marker genes (Pullin and McCarthy,, 2024).
@@
However, such a two-step pipeline for post-clustering differential analysis requires using the same data twice: first to estimate the clusters, and then again to estimate the differences between them and perform significance testing – a procedure sometimes referred to as “double-dipping” (Kriegeskorte et al.,, 2009). In post-clustering differential analysis, it has been shown that the primary risk of double dipping is to compromise the control of the Type I error rate of otherwise well-calibrated testing procedure for traditional differential analysis, leading to false positives (Zhang et al.,, 2019). Fundamentally, uncertainties stemming from cluster analysis outcomes, particularly related with determining the optimal number of clusters, could create artificial differences between homogeneous groups of observations. Failure to account for the double use of the data during the analysis may lead differential analysis tools to erroneously identify those artificial differences (Hivert et al., 2024a, ). Although challenges related to double dipping are increasingly studied, they remain unresolved in the context of scRNA-seq data analysis (Lähnemann et al.,, 2020). We described below the proposed approaches and the remaining challenges.
Leveraging the selective inference framework (Fithian et al.,, 2014), which involves conditioning on the clustering event during the derivation of the test statistic and the associated p-value, various methodological efforts have been proposed to address this double dipping issue in post-clustering inference (Gao et al.,, 2022). Zhang et al., (2019) introduced the Truncated-Normal test (TN-test), a four-step post-clustering differential analysis procedure that involves splitting the data into two parts, followed by cluster analysis on the first part of the data. Subsequently, a support vector machine (SVM) classifier is applied to the clustered data to learn the clustering structure and predict the cluster labels on the remaining data. Finally, a differential analysis is conducted between the predicted clusters using a truncated-normal test. This truncation, on either side of the hyperplane fitted by the SVM, allows to correct for the double dipping issue. However, information loss may occur as an inherent consequence of the data-splitting process. Moreover, the method involves multiple steps, introducing complexity into the overall analytical framework, and Song et al., (2023) highlighted the poor performances of the TN-test in their numerical simulations. More recently, Hivert et al., 2024a and Chen and Gao, (2023) have introduced univariate selective tests tailored to detect mean differences between two (multivariate) clusters. These methods explicitly condition on the clustering event to derive their p-value, relying on the set of all perturbed data sets that would yield the same partition when subjected to the same clustering algorithm. Hivert et al., 2024a rely on a Monte Carlo approach to suit any clustering algorithm, which comes at the expense of extensive computational times. On the other hand, Chen and Gao, (2023) explicitly describe this set specifically for $k-$means and hierarchical clustering applied to the squared distance matrix. Bachoc et al., (2023) also introduced a selective test designed for convex clustering. All these proposed selective methods designed for post-clustering inference either require specific clustering algorithms or introduce new specific test statistics. This consequently increases their complexity and makes their application less straightforward in the context of scRNA-seq data analysis.
@@
Leiner et al., (2023) have drawn inspiration from data splitting (which effectively addresses overfitting issues in machine learning) to break free from the selective inference framework. They propose a method called “data fission”, wherein the information within each individual observation is split into two independent parts. The first part could be used for cluster analysis, and labeling observations on the second part simply by matching them to their first counterpart. Differential analysis could then be conducted on the remaining information, i.e. the second part. However, in the context of post-clustering differential analysis, it is imperative that the two parts be independent, as each analysis (cluster analysis and differential analysis) must be performed independently to effectively prevent double dipping and the associated inflation of Type I error. Data fission requires strong parametric distributional assumptions, with only the Gaussian and Poisson (Neufeld et al.,, 2024) distributions ensuring independence between the two fissioned parts. Expanding upon this concept, Neufeld et al., 2023a have generalized the fission process by introducing “data thinning”. Building on the same foundational idea, they not only succeed in developing a process capable of decomposing data into more than two parts, but also broaden the spectrum of distributions where independence between each part is provided. This includes the negative binomial distribution, widely used when modeling RNA-seq data.
@@
Data fission and data thinning have emerged as promising alternatives to selective inference for post-clustering inference due to their compatibility with various clustering methods and differential analysis tests. However, they present some inherent limitations that make them difficult to apply for post-clustering inference. First, these methods lack results and justification when applied to mixture distributions, which are commonly used to model data with a clustered structure (Macqueen,, 1967). Consequently, the absence of such results inherently assumes a global null hypothesis of no clusters in the data when applying data fission or data thinning. Additionally, these methods assume prior knowledge of parameters (e.g., variance for the Gaussian distribution or overdispersion for the negative binomial distribution). Although robust estimators for these parameters could theoretically ensure the validity of the method, this further adds complexity for clustered data where each cluster has a different parameter value. In the absence of knowledge about the data structure, specifically the clusters, the only justifiable estimator is the full-sample one (i.e. computed using all the observations regardless of their mixture component) that fails to correctly estimate the intra-component parameter value.
In this article, we demonstrate that these approaches are not practical for real-world applications. Performing cluster analysis inherently assumes that the data originate from mixture models, contradicting the parametric assumptions of data fission and data thinning, which cannot decompose such distributions. Even if it is possible to move beyond the framework of mixture models by modeling each observation as a realization from distinct random variables with their own parameter values, accurately estimating scale parameters remains challenging without knowledge of the mixture. Specifically, we establish a link between the bias in estimating the variance parameter in the Gaussian distribution and the expected Type I error rate of the two-sample $t$-test (Welch,, 1947), underscoring the critical importance of a robust estimator. We employ a non-parametric method for estimating local variance in the Gaussian setting to try to use proximity between observations as a proxy of the unknown underlying mixture. However the poor performances of this approach demonstrates that accurate variance estimation, and hence control of the Type I error rate, remains challenging without knowing the mixture, i.e. the clustering, prior to decomposition.
@@ 2 Methods
In the following, capital letters represent random variables with a probability distribution function denoted as $p$, $x_{i}$ denotes a set of $n$ realizations of $X$, and matrices and vectors are bolden.
@@ 2.1 Data fission and data thinning
Let $\boldsymbol{X}$ be a random variable with a known distribution. Both data fission (Leiner et al.,, 2023) and its generalization, data thinning (Neufeld et al., 2023a, ), aim to decompose the random variable $\boldsymbol{X}$ into two (or more in data thinning) new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. These new variables are designed to i) retain information from the original variable $\boldsymbol{X}$, and ii) be independent. Of note, data fission can generate pairs of $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ that are not independent, but her we will focus on the independent cases only. The balance between the amount of information from $\boldsymbol{X}$ kept in either $\boldsymbol{X}^{(1)}$ or $\boldsymbol{X}^{(2)}$ is tuned by a hyperparameter $\tau$. Such a decomposition can be performed for various probability distributions of $\boldsymbol{X}$. For data fission, Leiner et al., (2023) identified only two distributions, the Gaussian and the Poisson, that satisfy the independence requirement. In contrast, Neufeld et al., 2023a established a comprehensive decomposition that ensures independence for all convolution-closed distributions. This encompasses Gaussian, Poisson (Neufeld et al.,, 2024) and Negative Binomial (Neufeld et al., 2023b, ) distributions.
Distribution of $\boldsymbol{X}$ $\boldsymbol{\tau}$ Data fission Data thinning
@@
$\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{p\times p})$
$\boldsymbol{X}^{(1)}=\boldsymbol{X}+\tau\boldsymbol{Z}$
$\boldsymbol{X}^{(2)}=\boldsymbol{X}-\frac{1}{\tau}\boldsymbol{Z}$
@@
Decompositions for the three distributions into two independent random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are detailed in Table 1. For proofs of independence, please refer to the Section $1$ of the Supplementary Materials for the Gaussian case, and to Neufeld et al., (2024) and Neufeld et al., 2023b for the Poisson and negative binomial distribution, respectively. Both Gaussian and negative binomial data fissions/thinnings require knowledge of scale parameters (namely $\boldsymbol{\Sigma}$ or $\theta$) for practical feasibility, as highlighted in Table 1. Theoretical guarantees, and especially the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are based on using the true values of these parameters. Yet, in real-life applications these are unknown and need to be estimated, most likely from the data as described in Neufeld et al., 2023b .

### v3

# Practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

@@ Abstract
Post-clustering inference in single-cell RNA sequencing (scRNA-seq) analysis presents significant challenges in controlling Type I error during differential expression analysis. Data fission, a promising approach that aims to split data into two independent parts, relies on strong parametric assumptions of non-mixture distributions that are inherently violated in clustered data. To address this limitation, we introduce conditional data fission, an extension designed to decompose each mixture component into two independent parts. However, we demonstrate that applying such conditional data fission to mixture distributions requires prior knowledge of the clustering structure to ensure valid post-clustering inference. This arises from the need to accurately estimate component-specific scale parameters, which are critical for performing decomposition while maintaining independence. We theoretically quantify how biases in estimating these parameters lead to inflated Type I error rates due to deviations from independence. Given that mixture components are typically unknown in practice, our results underscore the fundamental difficulty of applying data fission in real-world settings, despite its prior proposal as a solution for post-clustering inference.
Keywords: Double-dipping, Mixture model, Post-clustering inference, Type I error, Unsupervised learning
@@ 1 Introduction
Clustering encompasses all unsupervised statistical methods that group observations into homogeneous and separated clusters. Widely used in various application fields such as biology or genomics (Jaeger & Banks 2023), clustering plays a major role in revealing or summarizing the signals contained in different types of multivariate data. In single-cell RNA-seq (scRNA-seq) analysis, a high-throughput technique for measuring gene expression at the single-cell level, clustering is typically the first step used to explore cellular heterogeneity and functional diversity (Amezquita et al. 2020). This clustering groups cells according to their gene expression. Afterwards, the differential expression analysis (comparing gene abundance between groups) allows to identify and annotate cellular sub-populations. This leads to the identification of marker genes that could potentially serve as cell-type marker genes (Pullin & McCarthy 2024).
@@
However, such a two-step pipeline for post-clustering differential analysis requires using the same data twice: first to estimate the clusters, and then again to estimate the differences between them and perform significance testing – a procedure sometimes referred to as “double-dipping” (Kriegeskorte et al. 2009). In post-clustering differential analysis, it has been shown that the primary risk of double-dipping is to compromise the control of the Type I error rate of otherwise well-calibrated testing procedures for traditional differential analysis, leading to false positives (Zhang et al. 2019). Fundamentally, uncertainties stemming from cluster analysis results, particularly related to the determination of the optimal number of clusters, could create artificial differences between homogeneous groups of observations. Failure to account for the double use of data during the analysis may lead to erroneously identify those artificial differences as significant (Hivert et al. 2024). Although challenges related to double-dipping are increasingly studied, they remain unresolved in the context of scRNA-seq data analysis (Lähnemann et al. 2020).
@@
To address the double-dipping issue in post-clustering inference, various methodological developments have leveraged the selective inference framework (Fithian et al. 2017), which entails conditioning on the clustering event when deriving test statistics and associated p-values (Gao et al. 2022). For a recent and comprehensive review of existing methods, see Enjalbert Courrech et al. (2025). As an alternative to the selective inference paradigm, Leiner et al. (2023) have drawn inspiration from data splitting (which effectively addresses overfitting issues in machine learning) to break free from the selective inference framework. They propose a method called “data fission”, where the information within each individual observation is split into two independent parts. In theory, the first part can be used for clustering, and then for labeling observations in the second part simply by identifying them with their first matching counterpart. Differential analysis can then be conducted on the remaining information (i.e. the second part). However, in the context of post-clustering differential analysis, it is imperative that the two parts be independent, as each analysis (cluster analysis and differential analysis) must be performed independently to effectively prevent double-dipping and the associated inflation of Type I error. For this, data fission requires strong parametric distributional assumptions with only the Gaussian and Poisson (Neufeld, Gao, Popp, Battle & Witten 2024) distributions ensuring independence between the two fissioned parts. Expanding on this concept, Neufeld, Dharamshi, Gao & Witten (2024) have generalized the fission process by introducing “data thinning”. Building on the same foundational idea, they not only succeed in developing a process capable of decomposing data into more than two parts, but also broaden the spectrum of distributions where independence between each part is provided. This includes the negative binomial distribution, widely used for modeling RNA-seq data.
@@
Data fission and data thinning have emerged as promising alternatives to selective inference for post-clustering inference due to their compatibility with various clustering methods and differential analysis tests. However, they present some inherent limitations that make them difficult to apply for practical post-clustering inference. First, these methods lack results and justification when applied to mixture distributions, which are commonly used to model data with a clustered structure (Macqueen 1967). Consequently, the absence of such results inherently assumes a global null hypothesis of no clusters in the data when applying data fission or data thinning. Additionally, these methods assume prior knowledge of parameters (e.g., variance for the Gaussian distribution or overdispersion for the negative binomial distribution). Although robust estimators for these parameters could theoretically ensure the validity of the method, this further adds complexity for clustered data where each cluster has a different parameter value. In the absence of knowledge about the data structure, specifically potential clusters, the only justifiable estimator is the full-sample one (i.e. computed using all the observations regardless of their mixture component) that fails to correctly estimate the intra-component parameter value.
In this article, we demonstrate that these approaches are not practical for real-world applications. Performing cluster analysis inherently assumes that the data originate from mixture models, contradicting the parametric assumptions of data fission and data thinning, which cannot decompose such distributions. Although one could apply data fission without explicitly accounting for mixture distributions (i.e., marginally to the class membership), we show that this approach leads to incorrect inference due to induced correlations between the two fissioned parts. These correlations arise because marginal data fission relies on an estimate of the overall scale parameter, which is biased relative to the intra-component (conditional) scale parameter. Specifically, we establish a direct link between bias in variance estimation within a Gaussian mixture model and inflation of the Type I error rate in the two-sample t-test (Welch 1947), highlighting the necessity of a robust variance estimator. To ensure independence and valid inference while accounting for mixture structures, the ideal strategy would be to perform decomposition at the component level, that is conditionally on class membership. However, in real-world post-clustering inference, this approach is infeasible as class memberships are unknown and must be estimated through clustering itself, leading to a circular problem.
@@ 2 Methods
In the following, capital letters represent random variables with a probability distribution function denoted as $f$, $x_{i}$ denotes a set of $n$ realizations of $X$, and matrices and vectors are in bold.
@@ 2.1 Data fission and data thinning
Let $\boldsymbol{X}$ be a random variable with a known distribution. Both data fission (Leiner et al. 2023) and its generalization, data thinning (Neufeld, Dharamshi, Gao & Witten 2024), aim to decompose the random variable $\boldsymbol{X}$ into two (or more in data thinning) new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. These new variables are designed to i) retain information from the original variable $\boldsymbol{X}$, and ii) be independent. Of note, data fission can generate pairs of $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ that are not independent. Here, we focus only on the independent cases. The balance between the amount of information from $\boldsymbol{X}$ kept in either $\boldsymbol{X}^{(1)}$ or $\boldsymbol{X}^{(2)}$ is tuned by a hyperparameter $\tau$. Such a decomposition can be performed for various probability distributions of $\boldsymbol{X}$. For data fission, Leiner et al. (2023) identified only two distributions – Gaussian and Poisson – that satisfy the independence requirement. In contrast, Neufeld, Dharamshi, Gao & Witten (2024) established a comprehensive decomposition that ensures independence for all convolution-closed distributions. This encompasses Gaussian, Poisson (Neufeld, Gao, Popp, Battle & Witten 2024) and negative binomial (Neufeld et al. 2023) distributions.
Distribution of $\boldsymbol{X}$ $\boldsymbol{\tau}$ Data fission Data thinning Scale parameter
@@
$\boldsymbol{W}\sim\mathcal{N}(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{p\times p})$
$\boldsymbol{X}^{(1)}=\boldsymbol{X}+\tau\boldsymbol{W}$
$\boldsymbol{X}^{(2)}=\boldsymbol{X}-\frac{1}{\tau}\boldsymbol{W}$
@@
The decompositions for the three distributions into two independent random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are detailed in Table 1. For proofs of independence, please refer to Section $1$ of the Supplementary Materials for the Gaussian case, and to Neufeld, Gao, Popp, Battle & Witten (2024) and Neufeld et al. (2023) for the Poisson and negative binomial distribution, respectively. Data fission or thinning of the Gaussian and negative binomial distributions both require knowledge of a scale parameter (namely $\boldsymbol{\Sigma}$ or $\theta$, respectively) for practical feasibility, as presented in Table 1. Theoretical guarantees, and especially the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, rely on using the true values of these parameters. Yet, in real-life applications these are unknown and need to be estimated, most likely from the data as described in Neufeld et al. (2023).
2.2 Data fission and thinning for mixture models
2.2.1 Mixture models
Neufeld, Gao, Popp, Battle & Witten (2024) and Neufeld et al. (2023) have introduced data thinning as a framework to address post-clustering inference. However, existing decomposition approaches are limited to probabilistic models that assume homogeneity and fail to account for potential heterogeneity arising from distinct clusters. In other words, these methods rely on a global null hypothesis of no cluster in the data. When true clusters exist, a more suitable modeling approach involves mixture models (Bouveyron et al. 2019), where each component of the mixture corresponds to a distinct cluster.
Finite mixture models provide a probabilistic framework for capturing heterogeneity in clustered data by assuming that each observation arises from one of several latent subpopulations. Formally, let $X_{1},\dots,X_{n}$ be independent observations, and introduce a latent categorical variable $Z_{i}\in\{1,\dots,G\}$ that indicates the class membership of the observation $X_{i}$. The probability that a given observation $X_{i}$ belongs to the cluster $g$ is given by the mixing proportions $\pi_{g}=\mathbb{P}(Z_{i}=g),\text{ such that }\sum_{g=1}^{G}\pi_{g}=1$. Given that $X_{i}$ belongs to the cluster $g$ (i.e. $Z_{i}=g$), its conditional distribution $\mathbb{P}(X_{i}\mid Z_{i}=g,\theta_{g})$ follows a cluster-specific density function $f_{g}(X_{i}\mid\theta_{g})$ , parameterized by $\theta_{g}$.
The marginal distribution of $X_{i}$, obtained by summing over all possible latent classes, is then: $f(X_{i}\mid\theta)=\sum_{g=1}^{G}\pi_{g}f_{g}(X_{i}\mid\theta_{g})$. Mixture models thus offer a principled probabilistic framework for modeling structured heterogeneity, enabling the derivation of data fission and data thinning decompositions in the presence of clustered data.
We will restrict our analysis to the Gaussian setting where the conditional (i.e. cluster-specific) density function is given by:
$\mathbb{P}\left(\boldsymbol{X}_{i}\mid Z_{i}=g,\boldsymbol{\theta}_{g}\right)=f_{g}(\boldsymbol{X}_{i}\mid\boldsymbol{\theta}_{g})=\frac{1}{(2\pi)^{p/2}|\boldsymbol{\Sigma}_{g}|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{X}_{i}-\boldsymbol{\mu}_{g})^{T}\boldsymbol{\Sigma}_{g}^{-1}(\mathbf{X}_{i}-\boldsymbol{\mu}_{g})\right)$
for $g=1,\dots,G$, where $G>1$ denotes the number of mixture components. Each component is parametrized by its own parameter $\boldsymbol{\theta}_{g}=\left(\boldsymbol{\mu}_{g},\boldsymbol{\Sigma}_{g}\right)$ where $\boldsymbol{\mu}_{g}=\mathbb{E}\left[\boldsymbol{X}_{i}\mid Z_{i}=g\right]\in\mathbb{R}^{p}$ denotes the conditional (i.e. cluster-specific) mean vector and $\boldsymbol{\Sigma}_{g}=\operatorname{\mathbb{V}ar}\left(\boldsymbol{X}_{i}\mid Z_{i}=g\right)\in\mathbb{R}^{p\times p}$ is the conditional (i.e. cluster-specific) variance. Much of the results apply to the negative binomial case, with the overdispersion parameter being analogous to the variance parameter in the Gaussian case.
2.2.2 Conditional (per-component) data fission vs. marginal (global) data fission for mixture models
Since conditionally on $Z_{i}=g$, the random variables $\boldsymbol{X}_{i}$ follow a Gaussian distribution, i.e. $\boldsymbol{X}_{i}\mid Z_{i}=g\sim\mathcal{N}\left(\boldsymbol{\mu}_{g},\boldsymbol{\Sigma}_{g}\right)$, it is possible to conditionally decompose mixture models using traditional data fission or data thinning. Let $\boldsymbol{W}_{1},\dots,\boldsymbol{W}_{n}$ be independent Gaussian random variables such that, for all $g=1,\dots,G$, $\left(\boldsymbol{W}_{i}\mid Z_{i}=g\right)\overset{\text{iid}}{\sim}\mathcal{N}\left(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{g}\right)$. The data fission process, as described in Table 1, yields the following decomposition of each component of the mixture:
$\left(\boldsymbol{X}^{(1)}_{i}\mid Z_{i}=g\right)=\Big{(}\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i}\mid Z_{i}=g\Big{)}\,\text{and}\,\left(\boldsymbol{X}^{(2)}_{i}\mid Z_{i}=g\right)=\Big{(}\boldsymbol{X}_{i}-\frac{1}{\tau}\boldsymbol{W}_{i}\mid Z_{i}=g\Big{)}.$
This fission decomposition ensures that, for all $g=1,\dots,G$, the random variables $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ are conditionally independent given $Z_{i}$: $\left(\boldsymbol{X}_{i}^{(1)}\perp\!\!\!\perp\boldsymbol{X}^{(2)}_{i}\right)\mid Z_{i}=g$.
Thus, each mixture component is decomposed into two independent Gaussian random variables. However, this conditional data fission does not guarantee the marginal independence between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$. In particular, we can show that (see Section $2.1$ of Supplementary Materials):
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)=\sum\limits_{g=1}^{G}\pi_{g}\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)^{T}$ (1)
Notably, this conditional independence property of the decomposition holds only when the conditional variances $\boldsymbol{\Sigma}_{g}$ are used to sample the $\boldsymbol{W}_{i}$ and perform the conditional decomposition. However, in the context of post-clustering inference, the class memberships $Z_{i}$ are unknown and must be estimated through clustering. Consequently, the true conditional variances $\boldsymbol{\Sigma}_{g}$ are inaccessible and obtaining reliable estimators for them is a challenge. The only available estimate, without any prior knowledge of $Z_{i}$, is the empirical variance $\widehat{\boldsymbol{\Sigma}}$ which serves as an estimate of the marginal variance $\boldsymbol{\Sigma}$. However, $\widehat{\boldsymbol{\Sigma}}$ is a biased estimate of $\boldsymbol{\Sigma}_{g}.$ It becomes impossible to guarantee the conditional independence between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}.$ Nevertheless, it is always possible to perform a marginal data fission using the marginal variance $\boldsymbol{\Sigma}$ (or it estimate):
$\boldsymbol{X}^{(1)}_{i}=\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i}\quad\text{and}\quad\boldsymbol{X}^{(2)}_{i}=\boldsymbol{X}_{i}-\frac{1}{\tau}\boldsymbol{W}_{i}\quad\text{with}\quad\boldsymbol{W}_{i}\overset{\text{iid}}{\sim}\mathcal{N}\left(\boldsymbol{0}_{p},\boldsymbol{\Sigma}\right).$
Such a marginal fission ensures that $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ are uncorrelated but at the cost of introducing conditional and hidden dependencies given by (see Section $2.2$ of Supplementary Materials):
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}=g\right)=\boldsymbol{\Sigma}_{g}-\boldsymbol{\Sigma}.$ (2)
Covariance results on conditional and marginal data fission are summarized in Supplementary Table 1.
3 Results